OpenStudy (anonymous):

A long time ago, in a galaxy far, far away, Lama Su takes Obi-Wan Kenobi on a tour of the clone factory in Tipoca City on Kamino. In order to obtain the biomass necessary to construct the clones, a gigantic vat is filled with 700 gallons of a solution of liquefied Hutt fat containing 80 pounds of granulated Wookie hair. Then, pure liquefied Hutt fat is pumped into the vat at a rate of 6 gallons per minute, while the well-stirred mixture is removed from the vat at a rate of 9 gallons per minute. After 3 hours of pumping pure Hutt fat into the vat, how many pounds of granulated Wookie hair are left in the vat?

OpenStudy (kc_kennylau):

Can somebody convert 80 pounds into gallons thank you

OpenStudy (anonymous):

w = 8.33 q where w = weight of water (lbs) q = volume of water (US gallons) so q=w/8.33 -> q=80/8.33=9.60 gallons